# A Theory of Quantum Jumps and Its Applications to Lyman and Balmer Series

### v3

@@ 3.3 Radial probability density
Having obtained the normalized spatial wavefunction $\eta_{n}(r)$, the radial probability density in a spherical shell volume element can be evaluated, and it is given by
@@
${\cal{P}}_{n}(r_{a})\equiv\frac{dP_{n}(r_{a})}{dr_{a}}=\frac{1}{\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{ap}+3}\ r_{a}^{2(\beta_{ap}+1)}\ e^{-(2\beta_{ap}/n^{2})r_{a}}\ ,$ (19)
[…]
@@ 3.4 Physical meaning of the solutions
Having obtained the temporal and spatial solutions to the governing equation, the physical meaning of these solutions is now discussed. According to Eq. (12), the temporal component $\chi_{n}(t)$ of the wavefunction $\phi_{A}(t,\mathbf{r})$ decays exponentialy in time, and the rate of this exponential decay depends on the value of the characteristic frequency $\Omega_{n}$ given by Eq. (11). Let the time $T_{n}=2\pi/\Omega_{n}$ be the duration time of a single quantum jump; the results presented below show that the duration time is very short but finite. By calculating this time for a considered quantum jump, and knowing the distance travelled by the electron during the jump, the electron speed can be determined for each jump.
The spatial solution and its normalization factor given by Eqs (15) and (17), respectively, were used to obtain the radial probability density for any quantum jump (see Eq. 19). The theory predicts that after an electron makes a quantum jump, the probablility density of finding it is very narrow and centered at $r=n^{2}a_{B}$, which is consistent with Bohr’s rule for quantum jumps. Experimental results [6-8] demonstrated that the electron remains in the excited state for periods ranging from a few tenths of a second to a few seconds before jumping again, and returns to its original state. The developed theory describes such physical situations, and its solutions preserve the coherence of the wavefunction. To achieve it, the solutions are restricted only to the EM frequencies corresponding directly to the quantum jumps (see Sec. 4). In other words, the solutions do not account for measurements because they destroy the coherence of the wavefunction.
@@ 4.1 Lyman series absorption
The developed theory is now applied to a hydrogen atom by considering absorption of EM radiation that causes the electron’s transitions from its initial, $n_{i}=1$, to its final, $n_{f}=2$, $3$ or $4$ energy levels. For this Lyman series absorption, the frequencies $\omega_{o}$ and $\Omega_{n}$ are calculated (see Eqs 5 and 11) by using
$\omega_{o}=2.065\cdot 10^{16}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (20)
@@
$\Omega_{n}=2.03\cdot 10^{8}\ \frac{\sqrt{\omega_{o}}}{n}\ ,$ (21)
where $n=n_{f}$. The characteristic frequency $\Omega_{n}$ can be used to estimate the duration time of quantum jumps $T_{n}=1/\Omega_{n}$. Knowing the distance traveled by the electron and the duration time, the electron’s speed $v_{e}=(n_{f}^{2}-n_{i}^{2})a_{B}/T_{n}$ can also be calculated.
Lyman series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
@@
Having obtained $\beta_{B}$, the wavefunction $\eta_{n}(r_{a})$ given by Eq. (15) is calculated for the transitions from $n_{1}=1$ to $n_{f}=2$, $n_{f}=3$ and $n_{f}=4$. The wavefunctions corresponding to these transitions are normalized (see Eq. 17) with its units being $a_{B}^{-3/2}$ (e.g., [4]), and plotted in Fig. 1. The presented results demonstrate that the maxima of the wavefunctions are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, which is in agreement with Bohr’s rules for quantum jumps. Moreover, the wavefunctions are symmetric about their corresponding maxima, and the shapes of the wavefunctions change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$.
The computed wavefunctions are then used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition by using Eq. (19). The obtained results are ploted in Fig. 2. These are the probability densities of finding the electron at its new location after the EM radiation required for the jump was absorbed. The theory predicts that the resulting radial probablility densities are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, as postulated by Bohr in his rules for quantum jumps (e.g., [4]). The shapes of these probability curves range from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$. Comparison of the results of Fig. 2 to those in Fig. 1 shows the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$.
The theory predicts the characteristic frequency ($\Omega_{n}\sim 10^{16}$ $s^{-1}$), the time-scale ($T_{n}\sim 10^{-16}s$) for the considered jumps, which are used to estimate the electron’s speed during the jumps; according to Table 1 the speeds are $v_{e}\sim 10^{5}$ $m/s$. Moreover, the theory also gives the radial probability density of finding the electron in the excited state. As the experiments [7-9] demonstrated, the electron spends from a few tenths of a second to a few seconds in the excited state before it jumps back to its original orbital.
@@ 4.2 Balmer series absorption
Balmer series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
@@
The wavefunction $\eta_{n}(r_{a})$ for the transitions from $n_{1}=2$ to $n_{f}=3$, $n_{f}=4$ and $n_{f}=5$ is computed using Eq. (15) and plotted in Fig. 3; the presented wavefunctions are normalized (see Eq. 17) and their units are $a_{B}^{-3/2}$ (e.g., [4]). The location of the maximum of each plotted wavefuction is centered at $r_{a}$ = $9$, $16$ and $25$, respectively, which is consistent with Bohr’s rules for quantum jumps. Just as for the wavefunctions for the Lyman series absorption (see Fig. 1), the wavefunctions for the Balmer series absorption are also symmetric about their maxima, and their shapes change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$. These may be the characteristic features of all the series absorptions in a hydrogen atom.
@@
Having obtained the normalized wavefunctions for the Balmer series absorption, Eq. (19) is used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition, and the results are plotted in Fig. 4. As shown, the probability densities demonstrate that the most probable location of the electron after its absorption of the required EM radiation is $r_{a}$ = $9$, $16$ and $25$, respectively, which is in agreement with Bohr’s rules for quantum jumps (e.g., [4]). The shapes of these probability curves change from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$, similar as the probability densities for the Lyman series absorption shown in Fig. 2. Moreover, comparing Figs 4 and 3, the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$ is clearly seen.
@@ 5 Experimental verification
According to the results presented in Tables 1 and 2, the duration time, $T_{n}$, for the electron’s transitions corresponding to the Lyman and Balmer series absorptions is finite and of the order of $10^{-16}$ $s$ and $10^{-15}$ $s$, respectively. The fact that the duration time is finite, not instantenous as originally postulated (e.g., [1-5]), is consistent with the recent experimental results [13]. However, the predicted values of $T_{n}$ for the hydrogen atom cannot be directly compared to the times of quantum jumps measured by [13]. The reason is that the measurements in [13] were performed on the so-called artificial atoms, which are much larger in size than atoms, but they have discrete energy leveles. The times $T_{n}$ given in Tables 1 and 2 are theoretical predictions based on the developed theory and they still remain to be verfied experimentally.
The measurements performed by three different teams [7-9] demonstrated that individual atoms, after absorbing EM radiation, stay in their excited states for periods ranging from a few tenths of a second to a few seconds before jumping again. Several other indepedent experiments [10-12] confirmed this range. The theoretical results presented in this paper give the radial probablility densities in spherical shell volume elements for the Lyman and Balmer series absorptions, and these probability densities may last for time periods determined by the experiments. It is proposed that the predicted radial probability densities for a single quantum jump can be verified by using a quantum microscope similar to that designed by researchers [44], who used it to measure the orbital structure of Stark states in an excited hydrogen atom. There are other methods to measure quantum phenomena [45-48], but some of them may not be suitable to observe single orbitals in a hydrogen atom [47].

### v4

@@ 3.4 Radial probability density
After finding the normalized spatial wavefunction $\eta_{n}(r)$, the radial probability density in a spherical shell volume element can be evaluated, and it is given by
@@
${\cal{P}}_{n}(r_{a})\equiv\frac{dP_{n}(r_{a})}{dr_{a}}=\frac{1}{\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{B}+3}\ r_{a}^{2(\beta_{B}+1)}\ e^{-(2\beta_{B}/n^{2})r_{a}}\ ,$ (21)
which gives $\int_{0}^{\infty}{\cal{P}}_{n}(r_{a})dr_{a}=1$. The obtained ${\cal{P}}_{n}(r_{a})$ gives the radial probability density for the electron being in this state after the transition.
@@ 3.5 Physical meaning of the solutions
Having obtained the temporal and spatial solutions to the NAE, the physical meaning of these solutions is now discussed. According to Eq. (14), the temporal component $\chi_{n}(t)$ of the wavefunction $\phi_{A}(t,\mathbf{r})$ decays exponentialy in time, and the rate of this exponential decay depends on the value of the characteristic frequency $\Omega_{n}$ given by Eq. (13). Let $T_{n}=2\pi/\Omega_{n}$ be the duration time of a single quantum jump, then, the results presented below show that this time is very short but finite. By calculating this time for a considered quantum jump, and knowing the distance travelled by the electron during the jump, the electron speed can be calculated for each jump.
The spatial solution and its normalization factor given by Eqs (17) and (19), respectively, were used to obtain the radial probability density for any quantum jump (see Eq. 21). The theory predicts that after an electron makes a quantum jump, the probablility density of finding it is very narrow and centered at $r=n^{2}a_{B}$, which is consistent with Bohr’s rule for quantum jumps. Experimental results [6-8] demonstrated that the electron remains in the excited state for periods ranging from a few tenths of a second to a few seconds before jumping again, and returns to its original state. The developed theory describes such physical situations, and its solutions preserve the coherence of the wavefunction. To achieve it, the solutions are restricted only to the EM frequencies corresponding directly to the quantum jumps (see Sec. 4). In other words, the solutions do not account for measurements because they destroy the coherence of the wavefunction; to retain this coherence, Minev et al. [13] employed a special procedure in their experiment.
@@ 4.1 Lyman series absorption
The developed theory is now applied to a hydrogen atom by considering absorption of EM radiation that causes the electron’s transitions from its initial, $n_{i}=1$, to its final, $n_{f}=2$, $3$ or $4$ energy levels. For this Lyman series absorption, the frequencies $\omega_{o}$ and $\Omega_{n}$ are calculated from Eqs (7) and (13), which give
$\omega_{o}=2.065\cdot 10^{16}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ {\rm[s^{-1}]}\ ,$ (22)
@@
$\Omega_{n}=2.03\cdot 10^{8}\ \frac{\sqrt{\omega_{o}}}{n}\ {\rm[s^{-1}]}\ ,$ (23)
where $n=n_{f}$. The characteristic frequency $\Omega_{n}$ can be used to estimate the duration time of quantum jumps $T_{n}=2\pi/\Omega_{n}$. Knowing the distance traveled by the electron and the duration time, the electron’s speed $v_{e}=(n_{f}^{2}-n_{i}^{2})a_{B}/T_{n}$ can also be calculated.
Lyman series $n=n_{f}$ $\omega_{o}$ $s^{-1}$] $\Omega_{n}$ $s^{-1}$] $T_{n}$ [s] $v_{e}$ $m\ s^{-1}$]
@@
Having obtained $\beta_{B}$, the wavefunction $\eta_{n}(r_{a})$ given by Eq. (17) is calculated for the transitions from $n_{1}=1$ to $n_{f}=2$, $n_{f}=3$ and $n_{f}=4$. The wavefunctions corresponding to these transitions are normalized (see Eq. 19) with its units being $a_{B}^{-3/2}$ (e.g., [4]), and plotted in Fig. 1. The presented results demonstrate that the maxima of the wavefunctions are centered at $r_{a}$ = $4$, 99 and $16$, respectively, which is in agreement with Bohr’s rule for quantum jumps. Moreover, the wavefunctions are symmetric about their corresponding maxima, and the shapes of the wavefunctions change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$.
The computed wavefunctions are then used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition by using Eq. (21). The obtained results are ploted in Fig. 2. These are the probability densities for the electron being at its new location after the EM radiation required for the jump was absorbed. The theory predicts that the resulting radial probablility densities are centered at $r_{a}$ = $4$, 99 and $16$, respectively, as postulated by Bohr for quantum jumps (e.g., [4]). The shapes of these probability curves range from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$. Comparison of the results of Fig. 2 to those in Fig. 1 shows the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$.
The theory predicts the characteristic frequency ($\Omega_{n}\sim 10^{16}$ $s^{-1}$), the time-scale ($T_{n}\sim 10^{-16}s$) for the considered jumps, which are used to estimate the electron’s speed during the jumps; according to Table 1, the speeds are $v_{e}\sim 10^{5}$ $m/s$. Moreover, the theory also gives the radial probability density of finding the electron in the excited state. As the experiments [7-9] demonstrated, the electron spends from a few tenths of a second to a few seconds in the excited state before it jumps back to its original orbital.
@@ 4.2 Balmer series absorption
Balmer series $n=n_{f}$ $\omega_{o}$ $s^{-1}$] $\Omega_{n}$ $s^{-1}$] $T_{n}$ [s] $v_{e}$ $m\ s^{-1}$]
@@
The wavefunction $\eta_{n}(r_{a})$ for the transitions from $n_{1}=2$ to $n_{f}=3$, $n_{f}=4$ and $n_{f}=5$ is computed using Eq. (17) and plotted in Fig. 3; the presented wavefunctions are normalized (see Eq. 19) and their units are $a_{B}^{-3/2}$ (e.g., [4]). The location of the maximum of each plotted wavefuction is centered at $r_{a}$ = $9$, 1616 and $25$, respectively, which is consistent with Bohr’s rule for quantum jumps. Just as for the wavefunctions for the Lyman series absorption (see Fig. 1), the wavefunctions for the Balmer series absorption are also symmetric about their maxima, and their shapes change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$. These may be the characteristic features of all the series absorptions in a hydrogen atom.
@@
Having obtained the normalized wavefunctions for the Balmer series absorption, Eq. (21) is used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition, and the results are plotted in Fig. 4. As shown, the probability densities demonstrate that the most probable location of the electron after its absorption of the required EM radiation is $r_{a}$ = $9$, 1616 and $25$, respectively, which is in agreement with Bohr’s rule for quantum jumps (e.g., [4]). The shapes of these probability curves change from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$, similar as the probability densities for the Lyman series absorption shown in Fig. 2. Moreover, comparing Figs 4 and 3, the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$ is clearly seen.
4.3 Discussion of the obtained results
The results presented for the Lyman and Balmer series absorption are obtained by using Eqs (7) and (8) to specify the values of $\omega_{o}$ and $\mathbf{k_{o}}$ . Then, the NAE, which is a deterministic equation, is used to describe the transition from the initial state $n=n_{i}$ to the final state $n=n_{f}$. The transition is causal because it is triggered by EM radiation with given $\omega_{o}$ and $\mathbf{k_{o}}$, and the theory predicts the physical characteristics of the resulting electron’s jump.
In the developed theory, no dependence on $\theta$ and $\phi$ is considered as shown by Eq. (15), which depends only on $r$. As a result, all initial states are s-orbitals ($1s$ and $2s$ for the Lyman and Balmer series, respectively) with $l=0$. However, the external EM radiation that is specified by $\omega_{o}$ and $\mathbf{k_{o}}$ carries the momentum $l=1$, which makes the angular momentum of the electron to be non-zero at its final state. Thus, in this model, the final state of $\alpha, Ly$-$\beta$ and Ly-$\gamma$ are $2p$, $3d$ and $4f$ states respectively; none of these states has nodes in their radial probability densities as shown in Fig. 2.
Specifically, for Ly-$\alpha$ – one photon from $n=1$ to $n=2$ gives $l=1$; however, for Ly-$\beta$ – two photons from $n=1$ to $n=2$ and from $n=2$ to $n=3$ give $l=2$; and similar for Ly-$\gamma$. The reverse process (emission) is also described by the NAE, and it is assumed that in this process the electron returns to its original state immediately after the jump, or after some period of time that can be established experimetally (see Sect. 5 for discussion). Similar analysis can be done for the Balmer series absorption, and Fig. 4 shows that none of the resulting states has nodes in their radial probability densities.
Since quantum jumps are non-unitary processes, they cannot be described by the SE, which accounts only for unitary processes, thus, the NAE is needed. However, the SE can be used to compute the probability of finding possible states of a quantum system before any interaction with its surroundings takes place. Therefore, the fact that the initial and finals states for the considered Lyman and Balmer series are in agreement with those already known from the solutions to the SE is expected as these are the only states available for the electron in the hydrogen atom. With the SE giving the initial and final states, and $\omega_{o}$ and $\mathbf{k_{o}}$ corresponding to these states, the NAE describes the resulting quantum jumps, and gives the duration time for the jumps as well as it predicts the probability distribution in the final state, which depends on the initial state. Note that the final states would be different for different initial states, and that the NAE correctly predicts them. As this discussion shows, both the SE and NAE are needed to describe quantum jumps.
The modern version of Copenhagen interpretation requires that quantum jumps are random and abrupt, and that they obey Bohr’s rule (e.g., [3-5]). The results obtained in this paper based on the NAE are consistent with the Copenhagen interpretation as the NAE directly displays the Bohr rule, and the considered quantum jumps are random. However, the difference is that while in the Copenhagen interpretation quantum jumps are instantaneous, the NAE predicts their duration times to be very short but finite, and allows calculating such times, which can be verifiable by experiments.
According to the results presented in Tables 1 and 2, the duration time, $T_{n}$, for the electron’s transitions corresponding to the Lyman and Balmer series absorptions is finite and of the order of $10^{-16}$ $s$ and $10^{-15}$ $s$, respectively. Thus, the computed duration times for the considered quantum jumps are finite, not instantenous as originally postulated (e.g., [1-5]); their comparison to experimental data is presented and discussed in Sect. 5. The velocities of the quantum transitions given in Tables 1 and 2 are higher for the higher frequencies $\omega_{o}$ (or energies of EM radiation). These are expected results and they imply that the highest velocity would be reached for the electron that leaves the hydrogen atom during its ionization - see again Sect. 5 for more details.
As briefly described in the Introduction, quantum trajectory theory (QTT) based on the stochastic Schrödinger equation (e.g., [20,21,23]) allows finding non-real and subjectively real quantum trajectories of individual particles that obey the probabilities computed from the SE; this makes QTT compatible with the SE. There are some similarities between QTT and the NAE as both describe individual elementary particles. Thus, a modified version of QTT based on the NAE can be developed within the framework of subjective real quantum trajectories that require a detector to observe a quantum system, and the NAE’s parameters $\omega_{o}$ and $\mathbf{k_{o}}$ may play the role of such a ’detector’; however, such a task is out of the scope of this paper. Then, QTT and its modifed version would be complementary to each other, similarly to the SE and NAE.
An interesting result is that the periods corresponding to $\omega_{o}$, and the predicted values of the duration times $T_{n}$ for the Lyman and Balmer series given in Tables 1 and 2, respectively, are of similar orders. Nevertheless, the obtained theoretical results are consistent with the time-energy uncertainty relation $\Delta t\ \Delta E\geq\hbar/2$, which still remains a controversial issue in QM (e.g., [38], and references therein). Taking $\Delta t\sim T_{n}$ and $\Delta E=\hbar\omega_{o}$, the relation becomes $T_{n}\ \omega_{o}\geq 1/2$, and as the results of Tables 1 and 2 demonstrate, it is satisfied for all the theoretically predicted values. It must be noted that the predicted values of the duration times $T_{n}$ are very short, and their experimental verification is now considered.
@@ 5 Experimental verification
Measurements of quantum jumps are difficult because it requires a time-measuring device that captures the begining of a quantum jump. As a result, in the previous experiments [7-12], it was only confirmed that quantum jumps occured randomly and abruptly. In the experiment performed by Minev et al. [13], the time-scales associated with quantum jumps are of an order of microseconds. However, these measured time-scales cannot be directly compared to the duration times of quantum jumps predicted in this paper. The reason is that the measurements in [13] were performed with artificial atoms, which are much larger in size than atoms, but they have discrete energy levels; was suggested that the measured time-scales could be explained by quantum trajectory theory (QTT), but it was not explicitly demonstrated.
As the results of Tables 1 and 2 show, $T_{n}$ computed in this paper for the hydrogen atom are about $10$ orders of magnitude shorter. However, using Eqs (7) and (13), the frequencies $\omega_{o}$ and $\Omega_{n}$ can be calculated for a Rydberg atom with large values of $n$. In extreme cases, Rydberg states corresponding to $n>500$ or even $n>1000$ can also be considered (e.g., [41,42], and references therein). Simple estimates based on the presented theory demonstrate that in order to obtain $T_{n}\sim 10^{-6}$ $s$, it would require Rydberg states with $n>1000$. However, it remains to be determined whether the prediction made by using such extreme Rydberg states are relevant to the experimental data reported in [13].
The measurements of quantum jumps in shelved-electron experiments (e.g., [7-9], also [43] and references therein) demonstrated that individual atoms, after absorbing EM radiation, stay in their excited states for periods ranging from a few tenths of a second to a few seconds before jumping again; see also [44] for most recent results in artificial atoms. In such experiments, an electron is temporarily placed in a long-lived, shelved (non-fluorescent) state that turns off the atom’s fluorescence signal; the precise moment when the electron transitions back to a fluorescing state is identified as a quantum jump. Several other indepedent experiments (e.g., [10-13]) confirmed this range. The theoretical results presented in this paper do not describe this kind of experiments; however, it is suggested that the NAE could be used to predict radial probablility densities in spherical shell volume elements for the electron in its shelved (non-radiating) state.
Significant progress in absolute timing quantum jumps has been made since attosecond spectroscopy was applied to photoionization (e.g., [45] and references therein). Chronoscope measurements of the times involved in the photoelectric effect resulted in the duration of the primary photoexcitation process to be of the order of $10^{-17}$ $s$ [46,47]. On the other hand, estimates based on the Franck-Condon principle give the duration times for quantum jumps to be of the order of $10^{-15}$ $s$ [48]. A specific prediction based on the assumption that quantum jumps occur as a result of a resonance of the atomic electron with with the modes of the zero-point radiation field of Compton’s frequency, gives $5\cdot 10^{-17}$ $s$ [46]. The duration times for different quantum jumps corresponding to the Lyman and Balmer series given in Tables 1 and 2 lie well within the range of the empirical and theoretical evaluations reported in [45-48].
Measurements performed on a neutral helium atom, which was hit by a high energy laser pulse, showed that one of the electrons was ripped out of the atom causing helium to be ionized. The reported experimental results demonstrated that the process occured on time-scale of attoseconds (e.g., [45-47]). For the hydrogen atom considered in this paper, the frequency of EM radiation required for the atom’s ionization is $\omega_{o}=2.06\cdot 10^{16}$ $s^{-1}$. In the developed theory based on the NAE, the electron’s final state must be specified. Taking $n=n_{f}\sim 1000$, Eq. (13) gives the duration time to be of an order of attosecond. However, for the fully ionized atom $n=n_{f}\rightarrow\infty$, which is not allowed by the theory that is only valid if the final states of the electron are finite.
The theoretical results presented in this paper give the radial probablility densities in spherical shell volume elements for the Lyman and Balmer series absorptions. According to the experimental results reported by the NIST Physical Measurement Laboratory [49], the atomic lifetimes of $\alpha, Ly$-$\beta$ and Ly-$\gamma$ are $4.7\cdot 10^{-8}$ $s$, $0.56\cdot 10^{-8}$ $s$ and $0.13\cdot 10^{-8}$ $s$, respectively. However, for $H\alpha$, $H\beta$ and $H\gamma$, the same source gives $0.44\cdot 10^{-8}$ $s$, $0.084\cdot 10^{-8}$ $s$ and $0.025\cdot 10^{-8}$ $s$, respectively. Comparison of these atomic lifetimes to the results of Tables 1 and 2, it is seen that that the duration times $T_{n}$ are eight or nine orders of magnitude shorter than the the corresponding lifetimes. Therefore, it is suggested that the computed probabilty densities may be directly observed by using a quantum microscope similar to that designed by researchers [50] who used it to measure the orbital structure of Stark states in an excited hydrogen atom. There are other methods to measure quantum phenomena [51-53], but some of them may not be suitable to observe single orbitals in a hydrogen atom [53].
There are three main experimental limitations that might affect the detection of quantum jumps as predicted by the NAE. First, the uncertainty principle would affect the parameters $\omega_{o}$ and $\mathbf{k_{o}}$, and this uncertainty would affect the calculations of time scales and probability densities, introducing a degree of fuzziness. Second, measurement backaction caused by EM radiation used in a quantum microscope would alter the electron state, which is not fully captured by the NAE. Third, apparatus limitations (e.g., resolving power) lead to uncertainties in the determination of the electron’s position, which would affect the accuracy of the probability densities obtained from the NAE.
Having demonstrated that the NAE gives a solution to the quantum measurement problem [36], and that it alllows finding the duration times for quantum jumps as well as their resulting probability densities, as this paper shows, the NAE may also be used to develop new quantum-based technologies or improve/modify the currently known technologies used in cryptographic systems and in secure direct communications (e.g., [54]).